# Limits of Rauzy graphs of languages with subexponential complexity

### v1

@@ 2.2 Asymptotics of complexity function and convergence of Rauzy graphs
If ${\mathcal{L}}$ is prolongable, then $\alpha=0$ and this is [11, Proposition 2] of Cassaigne. He also treated the case $e_{l}(n)\equiv 1$ and $e_{r}(n)\equiv 0$ by using an auxiliary language ${\mathcal{L}}^{\prime}$ on $k+1$ letters. We will now adapt Cassaigne’s trick to the general case. Let $z$ be a new letter not in ${\mathcal{A}}$ and let ${\mathcal{A}}^{\prime}\coloneqq{\mathcal{A}}\cup\{z\}$. Define a new language ${\mathcal{L}}^{\prime}\subset{\mathcal{A}}^{\prime}$ by
@@
with $1\leq\max(\alpha+1,\beta)\leq\frac{3}{2}$. We can hence apply [11, Proposition 2] to ${\mathcal{L}}^{\prime}$ to obtain
@@ 2.3 Languages of one-sided subshifts
In this subsection, we finish the proof of Corollaries 1.2 and 1.3. The language of finite subwords of a one-sided subshift $\Sigma$ is always factorial and right-prolongable, but not necessarily prolongable. Hence, the only missing part is to prove that the conditions that we impose on a subshift imply almost prolongability of its language.
@@
The proof of Corollary 1.2 and Corollary 1.3 is now complete.
@@ 3 Limit of labelled Rauzy digraphs
As we already know, if the Benjamini-Schramm limit of (unlabelled) Rauzy digraphs for a subshift of unbounded subexponential complexity exists then it is $\vec{\mathbf{Z}}$. Throughout this section, we will assume that such a limit does exist. Now let us introduce labels on our Rauzy digraphs. Recall that we label an arc connecting two words $u$ and $v$ by the last letter of $v$. In this case, the limit of labelled Rauzy digraphs, if it exists, will be a probability measure on rooted labelled digraphs supported on $\vec{\mathbf{Z}}$ and invariant under the change of the root. In other words, the limit will be a probability measure on the set of bi-infinite words invariant under the shift. In this section, we show that if the subshift is uniquely ergodic then the Benjamini-Schramm limit of labelled Rauzy digraphs exists and is the unique ergodic invariant measure for the subshift. In the case of subshifts of linear complexity, this was earlier proved in [16].
@@
$\lim_{n}\sup_{k}|\frac{\lvert\omega_{k}\cdots\omega_{k+n}\rvert_{u}}{n+1}-F_{u}(\omega)|=0.$
@@
Let $u$ be a finite word of even length. The Benjamini–Schramm convergence of the Rauzy digraphs implies that there exists an unbounded non-decreasing sequence $\{d_{n}\}$ of integer numbers such that for any sufficiently large $n$, there exists a subset $E_{n}$ of vertices of ${\vec{\operatorname{Rauz}}}(n)$ such that $|E_{n}|=o(p(n))$ and each connected component of ${\vec{\operatorname{Rauz}}}(n)\setminus E_{n}$ is an oriented segment of length at least $d_{n}$.
For each segment of length $k$, we assign a finite sequence of labels that can be read following the orientation starting from the vertex with no ingoing edges. This sequence forms a finite subword of $\omega$ of length $k\geq d_{n}$. Due to the uniform frequencies condition, the number of occurrences of $u$ in this word is $F_{u}(\omega)k+o(k)$. Each such occurrence represents a vertex in ${\vec{\operatorname{Rauz}}}(n)$ whose $\frac{|u|}{2}$–neighborhood is the oriented segment whose sequence of labels is exactly $u$. Hence, the total number of such vertices in ${\vec{\operatorname{Rauz}}}(n)$ is $F_{u}(\omega)p(n)+o(p(n))$. Therefore, the labelled Rauzy digraphs converge to a measure on ${\mathcal{A}}^{\mathbf{Z}}$ such that the cylinder set with base $u$ has measure $F_{u}(\omega)$. That is this limit coincides with the unique invariant measure on $(\tilde{\Sigma}_{\omega},S)$.
@@
First, we apply Lemma 2.5 to ${\mathcal{L}}_{\Sigma}$ and find a sequence $\{n_{k}^{\prime}\}$ such that the underlying unlabelled Rauzy digraphs ${\vec{\operatorname{Rauz}}}(n_{k})$ converge to $\vec{\mathbf{Z}}$. Then, due to the compactness argument, we can choose a subsequence $\{n_{k}\}\subset\{n_{k}^{\prime}\}$ such that the Benjamini–Schramm limit along $n_{k}$ of the labelled Rauzy digraphs exists. This limit is supported on the set of labelings of $\vec{\mathbf{Z}}$ and is invariant under the change of the root, hence, it is a stationary measure on ${\mathcal{A}}^{\mathbf{Z}}$. Clearly, the measure of any cylinder set with base $u$ is positive only if $u$ belongs to the language of $(\Sigma,S)$. Hence, the limiting measure is a shift-invariant measure supported on the natural extension of the system. ∎
@@ 4 Examples and applications
In this section we explain how various dynamical systems fit in the framework of our main results (Theorem  1.1, Corollaries  1.2 and  1.3 and Theorem 1.4).
@@ 4.1 Symbolic dynamical systems
Recall that a substitution is a mapping from an alphabet $A$ to the set of words $A^{*}$. It can be extended to a monoid morphism $A^{*}\to A^{*}$ and further to a morphism $\phi:A^{\mathbb{N}}\to A^{\mathbb{N}}$. A pure morphic sequence is a fixed point of such a morphism. A natural generalization of this notion is that of a morphic sequence, that is, the image of a pure morphic sequence under some coding (a morphism $A\rightarrow B$ between two alphabets).
@@
Next, we consider Toeplitz subshifts.
Theorem A in [26] claims that for any rational number $\frac{p}{q}$ and any function $f(n)=o(n^{\alpha})$ such that $nf^{\prime}(n)=o(n^{\alpha})$ there exists a Toeplitz sequence whose complexity satisfies the following condition:
@@ 4.2 Interval Translation Mappings
Interval translation mappings (ITMs) were introduced by M. Boshernitzan and I. Kornfeld (see [8]) as a natural generalization of one of the most famous one-dimensional dynamical systems - interval exchange transformations (IETs). IETs were defined in 60s in connection with the study of measured foliations on oriented surfaces and billiards in rational polygons. Namely, IETs appear as the first return map to the transversal for the leaves of the measured foliation (or, equivalently, for the vertical translation flow on the translation surface or billiard flow in a rational billiard). We refer to [14, Chapter 5] for basics about IETs.
@@
Boshernitzan and Kornfeld ([8]) noticed that ITMs can be divided into two classes in the following way: looking at the sequence $\Sigma_{k}=I\cap T(I)\cdots\cap T^{k}(I)$ we distinguish two cases. Either there exists $m\in\mathbb{N}$ such that $\Sigma_{m}=\Sigma_{m+1}$, then we say that an ITM $T$ is of finite type. Morally it means that $T$ can be reduced to an IET. If such an $m$ does not exist and the limit of $\Sigma_{m}$ is a nowhere dense set, we say that $T$ is of infinite type. A challenging question is to understand the maps of infinite type because only these ITMs represent some dynamics crucially different from IETs. Additional motivation to study ITMs of infinite type comes from the geometric group theory, where a very similar object, named systems of isometries or band complex, appears in connection with actions of free groups on $\mathbb{R}$-trees (see [19] for the definitions); ITMs of infinite type can be seen as a particular case of systems of isometries of so called thin type.
@@ 4.3 Polygonal and polyhedral billiards
An important class of dynamical systems that fit in the framework of Theorem  1.2 but is not covered by Drummond’s results are billiards in rational polygons and, more generally, polyhedra.
@@ 5 A non uniquely ergodic example and subsequential limits of Rauzy digraphs
We conclude the paper with an example of a non-uniquely ergodic system where we identify a continuum of invariant measures with subsequential limits of labelled Rauzy digraphs. In a recent paper [12], Cassaigne and Kaboré constructed a uniformly recurrent infinite word $u$ with complexity bounded by $13n+$ 1 that does not have uniform frequencies; see also [25] for an older example with similar properties. The linear bound on the complexity implies that the unlabelled Rauzy digraphs converge to $\vec{\mathbf{Z}}$ due to Corollary 1.3. Their construction of such a word goes as follows. For certain very fast growing sequences $(l_{i})_{i}$, $(m_{i})_{i}$ and $(n_{i})_{i}$, consider two sequences $(u_{i})_{i}$ and $(v_{i})_{i}$ of finite binary words defined by the rule $u_{0}=0$, $v_{0}=1$ and for any non-negative integer $i$, $u_{i+1}=u_{i}^{m_{i}}v_{i}^{l_{i}}$ and $v_{i+1}=u_{i}^{m_{i}}v_{i}^{n_{i}}$. The word $u$ is the limit of finite words $u_{i}$’s. The exact values of the parameters that can be used for our purpose are given by

### v2

@@ 2.2 Asymptotics of complexity function and convergence of Rauzy graphs
If ${\mathcal{L}}$ is prolongable, then $\alpha=0$ and this is [13, Proposition 2] of Cassaigne. He also treated the case $e_{l}(n)\equiv 1$ and $e_{r}(n)\equiv 0$ by using an auxiliary language ${\mathcal{L}}^{\prime}$ on $k+1$ letters. We will now adapt Cassaigne’s trick to the general case. Let $z$ be a new letter not in ${\mathcal{A}}$ and let ${\mathcal{A}}^{\prime}\coloneqq{\mathcal{A}}\cup\{z\}$. Define a new language ${\mathcal{L}}^{\prime}\subset{\mathcal{A}}^{\prime}$ by
@@
with $1\leq\max(\alpha+1,\beta)\leq\frac{3}{2}$. We can hence apply [13, Proposition 2] to ${\mathcal{L}}^{\prime}$ to obtain
@@ 2.3 Languages of one-sided subshifts
In this subsection, we finish the proof of Corollary 1.2 and more generally show that Theorem 1.1 applies to one-sided subshifts possessing a finite subset with a dense orbit. The language of finite subwords of a one-sided subshift $\Sigma$ is always factorial and right-prolongable, but not necessarily prolongable. Hence, the only missing part is to prove that the conditions that we impose on a subshift imply almost prolongability of its language.
@@
The proof of Corollary 1.2 is now complete.
@@ 3 Limit of labelled Rauzy digraphs
As we already know, if the Benjamini-Schramm limit of (unlabelled) Rauzy digraphs for a subshift of unbounded subexponential complexity exists then it is $\vec{\mathbf{Z}}$. Throughout this section, we will assume that such a limit does exist. Now let us introduce labels on our Rauzy digraphs. Recall that we label an arc connecting two words $u$ and $v$ by the last letter of $v$. In this case, the limit of labelled Rauzy digraphs, if it exists, will be a probability measure on rooted labelled digraphs supported on $\vec{\mathbf{Z}}$ and unimodular. The limit can then be identified with a probability measure on the set of bi-infinite words invariant under the shift, as shown in Lemma 2.4. In this section, we show that if the subshift is uniquely ergodic then the Benjamini-Schramm limit of labelled Rauzy digraphs exists and is the unique ergodic invariant measure for the subshift. In the case of subshifts of linear complexity, this was earlier proved in [18].
@@
$\lim_{n}\sup_{k}\lvert\frac{\lvert\omega_{k}\cdots\omega_{k+n}\rvert_{u}}{n+1}-F_{u}(\omega)\rvert=0.$
@@
Let $u$ be a finite word of even length. The Benjamini–Schramm convergence of the Rauzy digraphs implies that there exists an unbounded non-decreasing sequence $\{d_{n}\}$ of integer numbers such that for any sufficiently large $n$, there exists a subset $E_{n}$ of vertices of ${\vec{\operatorname{Rauz}}}(n)$ such that $\lvert E_{n}\rvert=o(p(n))$ and each connected component of ${\vec{\operatorname{Rauz}}}(n)\setminus E_{n}$ is an oriented segment of length at least $d_{n}$.
For each segment of length $k$, we assign a finite sequence of labels that can be read following the orientation starting from the vertex with no ingoing edges. This sequence forms a finite subword of $\omega$ of length $k\geq d_{n}$. Due to the uniform frequencies condition, the number of occurrences of $u$ in this word is $F_{u}(\omega)k+o(k)$. Each such occurrence represents a vertex in ${\vec{\operatorname{Rauz}}}(n)$ whose $\frac{\lvert u\rvert}{2}$–neighborhood is the oriented segment whose sequence of labels is exactly $u$. Hence, the total number of such vertices in ${\vec{\operatorname{Rauz}}}(n)$ is $F_{u}(\omega)p(n)+o(p(n))$. Therefore, the labelled Rauzy digraphs converge to a measure on ${\mathcal{A}}^{\mathbf{Z}}$ such that the cylinder set with base $u$ has measure $F_{u}(\omega)$. That is, this limit coincides with the unique invariant measure on $(\tilde{\Sigma}_{\omega},S)$.
@@
First, we apply Lemma 2.7 to ${\mathcal{L}}_{\Sigma}$ and find a sequence $\{n_{k}^{\prime}\}$ such that the underlying unlabelled Rauzy digraphs ${\vec{\operatorname{Rauz}}}(n_{k})$ converge to $\vec{\mathbf{Z}}$. Then, due to the compactness argument, we can choose a subsequence $\{n_{k}\}\subset\{n_{k}^{\prime}\}$ such that the Benjamini–Schramm limit along $n_{k}$ of the labelled Rauzy digraphs exists. This limit is supported on the set of labelings of $\vec{\mathbf{Z}}$ and, by Lemma 2.4, can be identified with a shift-invariant measure on ${\mathcal{A}}^{\mathbf{Z}}$. Clearly, the measure of any cylinder set with base $u$ is positive only if $u$ belongs to the language of $(\Sigma,S)$. Hence, the limiting measure is a shift-invariant measure supported on the natural extension of the system. ∎
@@ 4 Examples and applications
In this section we explain how various dynamical systems fit in the framework of our main results (Theorem  1.1, Corollary 1.2 and Theorem 1.3).
@@ 4.1 Symbolic dynamical systems
Recall that a substitution is a mapping from an alphabet $A$ to the set of words $A^{*}$. It can be extended to a monoid morphism $A^{*}\to A^{*}$ and further to a morphism $\phi:A^{\mathbf{N}}\to A^{\mathbf{N}}$. A pure morphic sequence is a fixed point of such a morphism. A natural generalization of this notion is that of a morphic sequence, that is, the image of a pure morphic sequence under some coding (a morphism $A\rightarrow B$ between two alphabets).
@@
Next, we consider Toeplitz subshifts. Theorem A in [28] claims that for any rational number $\frac{p}{q}$ and any function $f(n)=o(n^{\alpha})$ such that $nf^{\prime}(n)=o(n^{\alpha})$ there exists a Toeplitz sequence whose complexity satisfies the following condition:
@@ 4.2 Interval Translation Mappings
Interval translation mappings (ITMs) were introduced by M. Boshernitzan and I. Kornfeld (see [10]) as a natural generalization of one of the most famous one-dimensional dynamical systems - interval exchange transformations (IETs). IETs were defined in 60s in connection with the study of measured foliations on oriented surfaces and billiards in rational polygons. Namely, IETs appear as the first return map to the transversal for the leaves of the measured foliation (or, equivalently, for the vertical translation flow on the translation surface or billiard flow in a rational billiard). We refer to [16, Chapter 5] for basics about IETs.
@@
Boshernitzan and Kornfeld ([10]) noticed that ITMs can be divided into two classes in the following way: looking at the sequence $\Sigma_{k}=I\cap T(I)\cdots\cap T^{k}(I)$ we distinguish two cases. Either there exists $m\in\mathbf{N}$ such that $\Sigma_{m}=\Sigma_{m+1}$, then we say that an ITM $T$ is of finite type. Morally it means that $T$ can be reduced to an IET. If such an $m$ does not exist and the limit of $\Sigma_{m}$ is a nowhere dense set, we say that $T$ is of infinite type. A challenging question is to understand the maps of infinite type because only these ITMs represent some dynamics crucially different from IETs. Additional motivation to study ITMs of infinite type comes from the geometric group theory, where a very similar object, named systems of isometries or band complex, appears in connection with actions of free groups on $\mathbf{R}$-trees (see [21] for the definitions); ITMs of infinite type can be seen as a particular case of systems of isometries of so called thin type.
@@ 4.3 Polygonal and polyhedral billiards
An important class of dynamical systems that fit in the framework of Theorem  1.1 but is not covered by Drummond’s results are billiards in rational polygons and, more generally, polyhedra.
@@ 5 A non uniquely ergodic example and subsequential limits of Rauzy digraphs
We conclude the paper with an example of a non-uniquely ergodic system where we identify a continuum of invariant measures with subsequential limits of labelled Rauzy digraphs. In a recent paper [14], Cassaigne and Kaboré constructed a uniformly recurrent infinite word $u$ with complexity bounded by $13n+$ 1 that does not have uniform frequencies; see also [27] for an older example with similar properties. The linear bound on the complexity implies that the unlabelled Rauzy digraphs converge to $\vec{\mathbf{Z}}$ due to Corollary 1.2.
Their construction of such a word goes as follows. For certain very fast growing sequences $(l_{i})_{i}$, $(m_{i})_{i}$ and $(n_{i})_{i}$, consider two sequences $(u_{i})_{i}$ and $(v_{i})_{i}$ of finite binary words defined by the rule $u_{0}=0$, $v_{0}=1$ and for any non-negative integer $i$, $u_{i+1}=u_{i}^{m_{i}}v_{i}^{l_{i}}$ and $v_{i+1}=u_{i}^{m_{i}}v_{i}^{n_{i}}$. The word $u$ is the limit of finite words $u_{i}$’s. The exact values of the parameters that can be used for our purpose are given by